OpenStudy (anonymous):

OfficeJet printer can copy Maria's dissertation in 14 min. The LaserJet printer can copy the same document in 22 min. If the two machines work together, how long would they take to copy the dissertation.

OpenStudy (anonymous):

we have to create and equation to solve this

OpenStudy (anonymous):

lets say: (1/14)+(1/22)=(1/x)

OpenStudy (anonymous):

if you can solve, this you will have the answer to your question :)

OpenStudy (anonymous):

I'm not seeing a common denominator there

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

of course there is an LCD

OpenStudy (anonymous):

maybe 14*22

OpenStudy (anonymous):

Ok yeah 308 so 22/308 + 14/308 = 36/308 = 18/154 = 9/77?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

hmm, check using product/sum: (14*22)/(14+22)=308/36=8.5555or8.6 hours

OpenStudy (anonymous):

ok yeah that makes sense, Thanks